The time, in minutes, taken to cook a chicken, C, is linearly related to the weight, in kilograms, of the chicken, k. It takes 40 minutes to cook a 1.4kg chicken and 50 minutes to cook a 1.8kg chicken.
Find the linear equation relating cooking time, C, with the weight of the chicken, k, for a 1.7kg Chicken

Answer by josgarithmetic(39617)

This is the same as a line with ordered pairs, (k,C) and two points are given; find the equation for the line. You are looking for C=mk+b, much like you were finding y=mx+b.

The points on this line are (1.4,40), and (1.8,50).
Slope, m=(50-40)/(1.8-1.4)=25.
From C=mk+b, b=C-mk. Pick a point and find b.
b=50-25*1.8=5

Equation would then seem, C=25k+5.

If k=1.7, C cooking time predicted is C=25*1.7+5=47.5 minutes